I have been given two numbers, 4,2/3 and I need to make this into a quadratic equation
Answer by ankor@dixie-net.com(22740)

given two numbers, 4, 2/3
From these we can derive two factors
x = 4
(x - 4)
and
x = 2/3
(3x - 2)
FOIL (x - 4)(3x - 2) = (3x^2 - 2x - 12x + 8) = (3x^2 - 14x + 8), is our equation